### v5

# Quantum State Preparation via Free Binary Decision Diagram

Abstract
[…]
I Introduction
Quantum state preparation (QSP) is the task of preparing a quantum state for a given classical description of the quantum state, which may be given by oracle access to the coefficients of the quantum state or by some data structure to store the classical description of the quantum state. QSP is a prerequisite as input for major quantum algorithms such as Harrow-Hassidim-Lloyd (HHL) algorithm Harrow et al. (2009), Hamiltonian simulation Low and Chuang (2019); Gilyén et al. (2019), and quantum machine learning with optimized random features Yamasaki et al. (2020); Yamasaki and Sonoda (2022); Yamasaki et al. (2023). QSP is also recognized as a potential bottleneck for quantum acceleration Aaronson (2015). Therefore, the problem of computational complexity of QSP is of great interest from both practical, and fundamental viewpoints.
In general, the classical description of an $n$-qubit quantum state may have $\exp(O(n))$ parameters, which are inherently hard to deal with in polynomial computational resources in $n$. In Refs. Bergholm et al. (2005); Plesch and Brukner (2011), it is shown that an $n$-qubit quantum state can be prepared by an $O(2^{n})$-depth quantum circuit composed of single-qubit unitary, and CNOT gates, without ancillary qubits. References Zhang et al. (2022); Rosenthal (2021); Sun et al. (2023) gave an $O(n)$-depth quantum circuit to prepare an $n$-qubit quantum state using single-, and two-qubit gates, but the cost of the required number of ancillary qubits grows exponentially in $n$. The trade-off between the circuit depth, and the number of ancillary qubits for QSP is studied in Ref. Rosenthal (2021).
[…]
However, in general, the QSP with $\mathrm{poly}(n)$-sized classical description of any $n$-qubit quantum state is known to be computationally hard Aharonov and Ta‐Shma (2007). Therefore, one challenge is to develop a quantum algorithm for QSP with an $O(\mathrm{poly}(n))$-sized quantum circuit to prepare a wider class of quantum states with $d=O(2^{n})$ nonzero coefficients.
[…]
blackTo address the challenge, we employ a data structure for representing Boolean functions, called binary decision diagrams (BDDs) Breitbar P et al. (1995). A $\mathrm{BDD}_{f}$ representing a Boolean function $f:x\in\{0,1\}^{n}\mapsto\{0,1\}$ is a rooted directed acyclic graph with one or two nodes of outdegree zero and other nodes of outdegree two, where each node in the BDD represents a Boolean variable. In this paper, our method specifically utilizes a free binary decision diagram (FBDD), rather than a binary tree (BT) or an OBDD used in standard BDD-based QSP methods. QSP using FBDDs has not been explored because FBDDs do not share a key property that both BTs and OBDDs have. The property is that the order of occurrences of the Boolean variables on each path from the root to the terminal nodes is consistent. It assures that, at each depth, all the Boolean variables corresponding to nodes at the same depth represent the same qubit. Leveraging this property, for example, the method in Ref. Prakash (2014) achieves quantum state preparation (QSP) by coherently accessing the qubits labeled by nodes at the same depth.
In this paper, we construct a quantum algorithm for QSP when the classical description of an $n$-qubit quantum state is given by an FBDD with weighted edges. We prove two claims. One is that any $\mathrm{FBDD}_{f}$ with $N$ nodes, and complex weighted edges represents a classical description of an $n$-qubit quantum state written by |$\ket{\psi_{f}}=\sum_{x\in\{0,1\}^{n}}c_{x}f(x)\ket{x}$, from which it is shown that the $\mathrm{FBDD}_{f}$ encodes a subset of the basis states designated by the Boolean function, i.e., the support {|$\{\ket{x}|\ x\in\{0,1\}^{n},f(x)=1\}$, and the weighted edges encode the coefficients $c_{x}$’s respectively. The other is that there exists a classical algorithm to construct a quantum circuit to prepare a quantum state represented by the weighted FBDD using $O(N)$ single-, and two-qubit gates, and $N$ ancillary qubits.
Depth # Ancillary qubits
# Amplitude
amplifications
Assumption
/ Oracle
An $n$-qubit quantum state
that can be prepared
FBDD-based
(This work)
$O({\rm poly}(n))$ $O({\rm poly}(n))$
Single-, and
two-qubit gates
Described by an FBDD
with ${\rm poly}(n)$ nodes.
OBDD-based Mozafari et al. (2020) $O(2^{n})$ 00 - Single-, and two-qubit gates
Described by an OBDD
with ${\rm poly}(n)$ nodes.
Mozafari et al. (2022) $O(kn)$ $1$
/ with $k$ paths.
Width-first Bergholm et al. (2005); Plesch and Brukner (2011) $O(2^{n})$ 00 -
Single qubit, and
CNOT gates
Generally applicable.
Depth-first Zhang et al. (2022); Rosenthal (2021) $O(n)$ $O(2^{n})$ Single-, and two-qubit gates Generally applicable.
Balanced Sun et al. (2023) $O(\frac{2^{n}}{m+n})$ $m\in\left[2n,O\left(\frac{2^{n}}{n\log n}\right)\right]$
Sparse state
preparation Zhang et al. (2022)
$O(\log(nd))$ $O(nd\log d)$
Single-, and
two-qubit gates
The number of nonzero
coefficients is $d$.
Grover-Rudolph Grover and Rudolph (2002) $O(n)$ $O(g)$
Coherent
arithmetic
Efficiently integrable
probability distribution.
Matrix product state
encoding Holmes and Matsuura (2020); García-Ripoll (2021)
$O(n)$ 00 -
Single-, and
two-qubit gates
Approximated by
$O(1)$ bond dimension.
Black-box Sanders et al. (2019) $O(g\log_{2}g)^{\ast 1}$ $12g+$ 1 $O(2^{n/2})^{\ast 2}$ Amplitude oracle Generally applicable.
Bausch (2022) $O(g)^{\ast 1}$ $g+\log_{2}g$
QET-based McArdle et al. (2022) $O\left(\frac{nd}{\mathcal{F}_{f}^{[N]}}\right)^{\ast 3}$ $4$
Single-, and
two-qubit gates
Polynomial or Fourier
series approximation.
Adiabatic state
preparation Wan and Kim (2020); Rattew and Koczor (2022)
$O\left(\frac{g^{2}}{\mathcal{F}^{4}_{f}}\right)^{\ast 4}$ $O(n+g)$
Amplitude
oracle
Efficiently computable
function.
[…]
When $N=\mathrm{poly}(n)$, our FBDD-based algorithm gives a $\mathrm{poly}(n)$-sized quantum circuit for QSP with $\mathrm{poly}(n)$ ancillary qubits, which reduces the ancillary qubits needed for existing QSP algorithms for preparing a general class of states Zhang et al. (2022); Sun et al. (2023); Rosenthal (2021). Our FBDD-based algorithm exponentially reduces the number of two-qubit gates used by a QSP algorithm Mozafari et al. (2020) based on ordered BDDs (OBDDs), a proper subset of FBDDs Breitbar P et al. (1995).
[…]
blackFurthermore, our FBDD-based algorithm also reduces exponentially the number of two-qubit gates used by another OBDD-based algorithm Mozafari et al. (2022) preparing an $n$-qubit quantum state described by an OBDD with $k=O(2^{n})$ paths. The approach in Ref. Mozafari et al. (2022) operates by preparing the paths in an OBDD one-by-one. As a result, the circuit complexity is linear in the number $k$ of paths in the OBDD, rather than in the number $N$ of nodes. It is important to emphasize that, even if $N$ is polynomial in the depth of the BDD, $k$ is not necessarily polynomial. In contrast, our method achieves circuit complexity that is linear in the number $N$ of nodes, which presents a significant advantage over the approach.
We give a non-trivial example of exponentially compressible data structures by using a symmetric Boolean function Breitbar P et al. (1995). This result shows the existence of $n$-qubit quantum states with $d=O(2^{n})$ nonzero coefficients that can be prepared by a $\mathrm{poly}(n)$-sized quantum circuit with $\mathrm{poly}(n)$ ancillary qubits using our FBDD-based algorithm, but requires exponential amount of ancillary qubits using the algorithm for QSP in Ref. Zhang et al. (2022).
[…]
blackReference Vinkhuijzen et al. (2024) provides an analytical investigation of three widely used classical descriptions of quantum states: MPS, decision diagrams (DDs), and restricted Boltzmann machines (RBMs). The authors map the relative succinctness of these data structures. Our results on weighted FBDDs are closely related to one of their results on the DDs called semiring-labeled decision diagrams (${\rm SLDD}_{\times}$), which is interpreted as weighted OBDDs in our paper’s terminology. This fact implies that the weighted FBDD can provide a more succinct classical description of quantum states than ${\rm SLDD}_{\times}$.
We discuss four possible applications. First, we explore to prepare a uniform superposition of basis states designated by a Boolean function $f:x\in\{0,1\}^{n}\mapsto f(x)\in\{0,1\}$. We define a quantum state encoding the Boolean function $f$ as |$\ket{\phi_{f}}:=|f|^{-\frac{1}{2}}\sum_{x\in\{0,1\}^{n}}f(x)\ket{x}$, i.e., the amplitude encoding of the Boolean function, where $|f|$ is the number of $x\in\{0,1\}^{n}$ such that $f(z)=1$. Then, for a given $\mathrm{FBDD}_{f}$ representing the Boolean function $f$, we propose a classical algorithm with $O(N)$ queries to the nodes in the ${\rm FBDD}_{f}$ assigning weights to the $\mathrm{FBDD}_{f}$ to derive a classical description of the quantum state encoding $f$.
Second, we give another encoding of the Boolean function $f$ represented by an ${\rm FBDD}_{f}$. We give a classical algorithm of constructing a quantum circuit mapping |$\ket{x}$ to $e^{i\theta f(x)}\ket{x}$ using $O(N)$ Toffoli gates, and $N-1$ ancillary qubits.
Third, we provide another example of an $n$-qubit quantum state that can be prepared by an $O(n^{2})$-sized quantum circuit using $O(n^{2})$ ancillary qubits. We show that the quantum state cannot be prepared efficiently by a black-box QSP algorithm Bausch (2022) with respect to the number of amplitude amplifications.
Forth, we discuss that our proposed techniques of QSP can be applied to the block encoding, a technique of encoding a matrix as a block of a unitary, defined in Ref. Gilyén et al. (2019). We give two non-trivial examples of the block encodings of non-sparse matrices that are efficiently implemented.
The rest of the paper is organized as follows. In Sec II, we introduce a binary decision diagram (BDD), and its variants, a free binary decision diagram (FBDD), and an ordered binary decision diagram (OBDD). Then, we define an FBDD with weighted edges. In Sec. III, we prove that any $\mathrm{FBDD}_{f}$ with complex weighted edges is a classical description of an $n$-qubit quantum state whose support is {|$\{\ket{x}|\ x\in\{0,1\}^{n},f(x)=1\}$, and propose a classical algorithm to construct a quantum circuit to prepare the quantum state. We give an example of a $O(\mathrm{poly}(n))$-sized $\mathrm{FBDD}_{f}$ that is a classical description of a quantum state with exponentially large nonzero coefficients. In Sec. IV, we show a classical algorithm that assigns weights to any $\mathrm{FBDD}_{f}$ to derive a classical description of a quantum state encoding $f$. In Sec. V, for a given ${\rm FBDD}_{f}$, we show a classical algorithm constructing a quantum circuit mapping |$\ket{x}$ to $e^{i\theta f(x)}\ket{x}$ directly. In Sec. VI, we provide a ${\rm WFBDD}_{f}$ using $n^{2}$ nodes. Then, we show that our FBDD-based algorithm for QSP can generate the quantum state represented by the ${\rm WFBDD}_{f}$, but a black-box QSP algorithm Bausch (2022) cannot efficiently with respect to the number of the amplitude amplifications. In Sec. VII, we discuss that the techniques of our FBDD-based QSP can be applied to the block encoding, and show two non-trivial examples of the block encodings of non-sparse matrices. Finally, we give a conclusion.
Figure 1: A free binary decision diagram ($\mathrm{FBDD}_{f}$) representing a Boolean function $f(x_{1},x_{2},x_{3},x_{4})=x_{1}x_{3}x_{4}+\bar{x}_{1}x_{2}x_{4}+\bar{x}_{1}\bar{x}_{2}\bar{x}_{3}+\bar{x}_{1}x_{2}\bar{x}_{3}\bar{x}_{4}$. The $\mathrm{FBDD}_{f}$ is a rooted directed acyclic graph with one or two (rectangle) nodes of outdegree zero, and the other (circled) nodes of outdegree two. The rectangle nodes are labeled with 00, and $1$, called the 00-terminal node, and the $1$-terminal node. The other circle nodes are labeled with the variables $x_{1},x_{2},x_{3}$, and $x_{4}$, called the internal nodes. One node of indegree zero is called the root. Two outgoing edges of each internal node are labeled with 00, and $1$ respectively, called the 00-edge, and the $1$-edge, and shown as dotted, and solid lines. A path from the root to one of the terminal nodes represents a pair of an input bit sequence, and the output bit. For given $x_{1},x_{2},x_{3},x_{4}\in\{0,1\}$, $f(x_{1},x_{2},x_{3},x_{4})$ can be evaluated by starting from the root, and selecting the $x_{i}$-edge in each node until reaching the terminal nodes. If there is no node labeled with a variable on the path, the assignment of the variable is arbitrary. For instance, $x_{1},x_{2},x_{3},x_{4}=0,0,0,0$, and $0,0,0,1$ share the same path. Note that the word ”free” means that the order of occurrences of the variables on each path from the root to the terminal nodes is not fixed.
II Preliminaries
In this section, we give some definitions used in this study. Let $b$, $[n]$, and $[n]_{0}$ denote the bit $\{0,1\}$, $\{1,\ldots,n\}$, and $\{0,\ldots,n-1\}$ respectively. First, we introduce a binary decision diagram (BDD), and its variants, a free binary decision diagram (FBDD), and an ordered binary decision diagram (OBDD). Then, we define an FBDD with weighted edges (WFBDD).
A binary decision diagram (BDD) is a rooted directed acyclic graph with one or two nodes of outdegree zero, and the other nodes of outdegree two, to represent a Boolean function $f(x_{1},\ldots,x_{n})$ of variables $x_{1},\ldots,x_{n}$. (See Fig. 1.) The nodes of outdegree zero are labeled with 00, and $1$ respectively, called the 00-terminal node or the $1$-terminal node. The other nodes of outdegree two are labeled with the variables $x_{1},\ldots,x_{n}$, called the internal nodes. There is one internal node of indegree zero, called the root. Two outgoing edges of each internal node are labeled with 00, and $1$ respectively, called the 00-edge, and the $1$-edge of the internal node. A path from the root to one of the terminal nodes represents a pair of an input bit sequence, and the output bit. For given $x_{1},\ldots,x_{n}\in\{0,1\}$, $f(x_{1},\ldots,x_{n})$ can be evaluated by starting from the root, and selecting the $x_{i}$-edge in each node until reaching the terminal nodes. If there is no node labeled with a variable on the path, the assignment of the variable is arbitrary.
Although a BDD has redundant nodes in general, we can reduce the number of redundant nodes in a given BDD according to two contraction rules, called the redundant node deletion, and the equivalent node sharing. The redundant node deletion is the deleting of an internal node when two outgoing edges from the internal node have the same head node. The equivalent node sharing is the merging of two internal nodes when the two internal nodes share both of the head nodes of the two outgoing edges from their respective internal nodes. A BDD to which the contraction rules are applied enough are called the reduced BDD (RBDD). (See Fig. 2.) In summary, we give a formal definition of RBDDs as follows.
Figure 2: Contraction rules, and examples. (a) The redundant node deletion is the deleting of an internal node when two outgoing edges from the internal node have the same head node, where the triangle represents a partial $\mathrm{BDD}_{F}$ representing a Boolean function $F$, and the top vertex of the triangle means the root of the $\mathrm{BDD}_{F}$. (b) The equivalent node sharing is the merging of two internal nodes when the two internal nodes share both of the head nodes of the two outgoing edges from their respective internal nodes. (c) A $\mathrm{BDD}_{f}$ representing a Boolean function $f(x_{1},x_{2},x_{3})=\bar{x}_{1}\bar{x}_{2}x_{3}+x_{1}\bar{x}_{2}x_{3}+x_{1}x_{2}\bar{x}_{3}$, and the application of the redundant node deletion: The internal node $u_{4}$ is deleted because the head nodes of the 00-edge, and the $1$-edge are the same 00-terminal node $t_{0}$. (d) The application of the equivalent node sharing, and the $\mathrm{RBDD}_{f}$. The internal node $u_{5}$ is merged to another internal node $u_{3}$ because both of the head nodes of the two outgoing edges from their respective internal nodes are the 00-internal node $t_{0}$, and the $1$-internal node $t_{1}$.
Definition 1.
A reduced binary decision diagram $\mathrm{RBDD}_{f}=(V,E)$ representing a Boolean function $f(x_{1},\ldots,x_{n})$ of variables $x_{1},\ldots,x_{n}$ is a rooted directed acyclic graph $G=(V,E)$ with one or two nodes of outdegree zero, the other nodes of outdegree two, and the following properties.
  • 1
    The nodes of outdegree zero are labeled with 00, and $1$ respectively, called the 00-terminal node $t_{0}$, and the $1$-terminal node $t_{1}$.
  • 2
    The other nodes of outdegree two are labeled with the variables $x_{1},\ldots,x_{n}$, called the internal nodes. There is one internal node of indegree zero, called the root $r$. Let $i(u)$ denote the index of the variable labeled to an internal node $u$.
  • 3
    Two outgoing edges of each internal node are labeled with 00, and $1$ respectively, called the 00-edge, and the $1$-edge of the internal node. Let $e_{b}(u)$, and $h_{b}(u)$ denote the $b$-edge of an internal node $u$, and the head node of the $b$-edge.
  • 4
    The redundant node deletion cannot be applied. For any internal node $u$, two head nodes of the two outgoing edges from $u$ are different, i.e., $h_{0}(u)\neq h_{1}(u)$.
  • 5
    The equivalent node sharing cannot be applied. Any two internal nodes $u,v$ labeled with the same variable do not share both of the head nodes of two outgoing edges from $u$, and $v$, i.e., when $i(u)=i(v)$, $(h_{0}(u),h_{1}(u))\neq(h_{0}(v),h_{1}(v))$.
From Ref. Breitbar P et al. (1995), there are several classes of binary decision diagrams (BDDs). In this study, to use the results of Ref. Breitbar P et al. (1995), we introduce the following free BDDs, and ordered BDDs. (See Fig. 3.)
Figure 3: FBDD, OBDD, and WFBDD for $f=\bar{x}_{1}\bar{x}_{2}x_{3}+\bar{x}_{1}x_{2}x_{4}+x_{1}\bar{x}_{3}x_{4}+x_{1}x_{2}x_{3}$. (a) In the FBDD, each path from the root to the terminal nodes has at most one occurrence of each variable. (b) In the OBDD, the order of occurrences of the variables on each path from the root to the terminal nodes is consistent. (c) In the WFBDD, directed edges are assigned weights. If the head node of a directed edge is the 00-terminal node, the weight of that edge is set to zero.
Definition 2.
A (reduced) free binary decision diagram $\mathrm{FBDD}_{f}$ representing a Boolean function $f(x_{1},\ldots,x_{n})$ of variables $x_{1},\ldots,x_{n}$ is an $\mathrm{RBDD}_{f}$ for which each path from the root to the terminal nodes has at most one occurrence of each variable.
Definition 3.
A (reduced) ordered binary decision diagram $\mathrm{OBDD}_{f}$ representing a Boolean function $f(x_{1},\ldots,x_{n})$ of variables $x_{1},\ldots,x_{n}$ is an $\mathrm{FBDD}_{f}$ such that the order of occurrences of the variables on each path from the root to the terminal nodes is consistent with some linear order.
For the quantum state preparation, we encode the support of a target state in the FBDD. On the other hand, we encode coefficients of the state into weights of the edges of the FBDD. To establish this data structure more formally, we define the FBDD with weighted edges, called the weighted FBDD. (See Fig. 3.)
Definition 4.
A weighted free binary decision diagram $\mathrm{WFBDD}_{f}=(V,E)$ is an $\mathrm{FBDD}_{f}$ representing a Boolean function $f(x_{1},\ldots,x_{n})$ of variables $x_{1},\ldots,x_{n}$ with the following properties.
  • 1
    Complex weights are assigned to directed edges. Let $w_{b}(u)$ denote the weight of the $b$-edge of an internal node $u$.
  • 2
    When the head node of any directed edge is the 00-terminal node, the weight assigned to the directed edge is zero, i.e., for any internal node $u$, when $h_{b}(u)=t_{0}$, $w_{b}(u)=0$.
This definition means that the WFBDD can be regarded as the BDD to which the contraction rules are applied first, and weights are assigned after the contraction. Note that a weighted $\mathrm{OBDD}$ ($\mathrm{WOBDD}$) is also defined by assigning nonzero weights to all directed edges of a given $\mathrm{OBDD}$ except ones of which the head nodes are the 00-terminal node.
III Quantum state preparation via weighted FBDD
In this section, we show a quantum algorithm for QSP via weighted FBDD. In Sec. III.1, we prove that any $\mathrm{FBDD}_{f}$ with complex weighted edges is a classical description of an $n$-qubit quantum state whose support is {|$\{\ket{x}|\ x\in\{0,1\}^{n},f(x)=1\}$. Then, we give an example of an $O(\mathrm{poly}(n))$-sized $\mathrm{FBDD}_{f}$ that is a classical description of a quantum state with exponentially large nonzero coefficients. In Sec. III.2, we show that any quantum state represented by the weighted FBDD with $N$ nodes can be prepared by an $O(N)$-sized quantum circuit using $N$ ancillary qubits.
III.1 Classical description of quantum states by weighted FBDD
In this study, we construct all the quantum circuits by single-qubit unitaries, the controllization of them, and Toffoli gates. In this section, we prove that an arbitrary $\mathrm{WFBDD}_{f}$ is a classical description of an $n$-qubit quantum state whose support is {|$\{\ket{x}|\ x\in\{0,1\}^{n},f(x)=1\}$. First, we define a quantum state described by a given $\mathrm{WFBDD}_{f}$ as follows.
Definition 5.
For a given $\mathrm{WFBDD}_{f}$, and $z\in\{0,1\}^{n}$, let $P(z)=(V_{z},E_{z})$ be the path derived by starting from the root, and selecting the $z_{i}$-edge in each internal node until reaching one of the terminal nodes, where $z_{i}$ is the $i$-th bit of $z$. Then, the $n$-qubit quantum state described by the $\mathrm{WFBDD}_{f}$ is defined by
|$\ket{\psi}=\sum_{z\in\{0,1\}^{n}}\alpha(z)\ket{z},$ (1)
where the amplitude $\alpha(z)$ is given by
$\alpha(z)=\frac{1}{\sqrt{2^{n-|V_{z}|+1}}}\prod_{u\in V_{z}\setminus\{t_{0},t_{1}\}}\frac{w_{z_{i(u)}}(u)}{\sqrt{|w_{0}(u)|^{2}+|w_{1}(u)|^{2}}}.$ (2)
Note that $\alpha(z)=0$ when $P(z)$ reaches the terminal node $t_{0}$ since the edge weight to $t_{0}$ is defined to be zero.
The path $P(z)$ specifies a sequence of nodes from the root to one of the terminal nodes. At each node $u$, the amplitude value is successively updated according to the ratio of two weights. The first multiplicative factor in Eq. (2) comes from the deleted nodes having equal weighted edges. Let us show the above definition is well-defined, i.e., $|\alpha(z)|^{2}$ is a probability distribution.
Lemma 1.
The square of the absolute value of the amplitude $\alpha(z)$ in Eq. (2) is a probability distribution, i.e., $|\alpha(z)|^{2}\in[0,1]$ for $z\in\{0,1\}^{n}$, and $\sum_{z}|\alpha(z)|^{2}=1$.
Proof.
It is trivial that $|\alpha(z)|^{2}\in[0,1]$ because of the definition of Eq. (2). Then, let us show the probability conservation. In this proof, the goal is to show that
$\sum_{z:f(z)=1}|\alpha(z)|^{2}=1,$ (3)
because $\alpha(z)=0$ for $z$ such that $f(z)=0$ from the definition of $\alpha(x)$. As preparation, denoting $N_{c}(u)$ as the head nodes of the directed edges out from an internal node $u$, let us define the layers as follows:
  • 1.
    $L_{1}=\{t_{0},t_{1}\}$;
  • 2.
    for $i>1$, $L_{i}=\{u\in V\setminus\{t_{0},t_{1}\}\ |\ N_{c}(u)\subseteq L_{1}\cup\cdots\cup L_{i-1}\ {\rm,and}\ N_{c}(u)\cap L_{i-1}\neq\emptyset\}$.
(See Fig. 4 as an example.)
Figure 4: An example of layers in the proof of Lemma 1. The first layer $L_{1}$ is $\{t_{0},t_{1}\}$ by definition. The second layer $L_{2}$ is $\{u_{4},u_{5}\}$ since all the head nodes of the directed edges out from $u_{4}$ and $u_{5}$ belong to $L_{1}$. The remaining layers $L_{3}$, $L_{4}$, and $L_{5}$ can be defined sequentially in this way.
Note that $V=\bigcup_{i}L_{i}$ because if there exists an internal node $u\not\in\bigcup_{i}L_{i}$, then at least one of the two head nodes of the directed edges out from $u$ is not in $\bigcup_{i}L_{i}$, and repeating this leads to the conclusion that either $t_{0}$ or $t_{1}$ is not in $L_{1}$. Furthermore, let $P_{u}=(V_{u},E_{u})$ be one path derived by starting from the root to the node $u$. Then, let us define $B_{P_{u}}\subseteq\{0,1\}^{n}$ as a set of $z$ such that all the paths from the root to the terminal node $t_{1}$ containing the path $P_{u}$.
First, we prove the following by induction. For any node $u$, and one path $P_{u}$,
$\sum_{z\in B_{P_{u}}}|\alpha(z)|^{2}=\left\{\begin{array}[]{ll}1,&u=r,\\ \prod_{v\in V_{u}\setminus\{u\}}\frac{|w_{z_{i(v)}}(v)|^{2}}{|w_{0}(v)|^{2}+|w_{1}(v)|^{2}},&u\neq r.\end{array}\right.$ (6)
For a given $u\in L_{1}$, noting the multiplicity of $z$ for deleted nodes on the path $P_{u}$, we conclude the claim from the definition of $\alpha(z)$. Then, let us assume that Eq. (6) is satisfied for $u\in L_{1}\cup\cdots\cup L_{i}$. For a given $u\in L_{i+1}$, consider two head nodes $h_{0}(u)$, and $h_{1}(u)$ of the 00-edge $e_{0}(u)$, and the $1$-edge $e_{1}(u)$, respectively. There are two cases of $h_{0}(u)$, and $h_{1}(u)$. One is the case that either of $h_{0}(u)$, and $h_{1}(u)$ is $t_{0}$. The other is that both of $h_{0}(u)$, and $h_{1}(u)$ are in $\{t_{1}\}\cup L_{2}\cup\cdots\cup L_{i}$. For the first case, assuming that $h_{0}(u)=t_{0}$ without loss of generality, $B_{P_{u}}=B_{P_{h_{1}(u)}}$. Then, we can derive that
$\sum_{z\in B_{P_{u}}}|\alpha(z)|^{2}=\sum_{z\in B_{P_{h_{1}(u)}}}|\alpha(z)|^{2}=\prod_{v\in V_{u}\setminus\{u\}}\frac{|w_{z_{i(v)}}(v)|^{2}}{|w_{0}(v)|^{2}+|w_{1}(v)|^{2}}\frac{|w_{1}(u)|^{2}}{|w_{1}(u)|^{2}}=\prod_{v\in V_{u}\setminus\{u\}}\frac{|w_{z_{i(v)}}(v)|^{2}}{|w_{0}(v)|^{2}+|w_{1}(v)|^{2}}.$ (7)
In the second case, since $B_{P_{u}}=B_{P_{h_{0}(u)}}\cup B_{P_{h_{1}(u)}}$, we can derive that
$\sum_{z\in B_{P_{u}}}|\alpha(z)|^{2}=\prod_{v\in V_{u}\setminus\{u\}}\frac{|w_{z_{i(v)}}(v)|^{2}}{|w_{0}(v)|^{2}+|w_{1}(v)|^{2}}\frac{|w_{0}(u)|^{2}+|w_{1}(u)|^{2}}{|w_{0}(u)|^{2}+|w_{1}(u)|^{2}}=\prod_{v\in V_{u}\setminus\{u\}}\frac{|w_{z_{i(v)}}(v)|^{2}}{|w_{0}(v)|^{2}+|w_{1}(v)|^{2}}.$ (8)
Therefore, Eq. (6) is satisfied. Finally, applying the above discussion on the root $r$, we can prove the claim. ∎
Note that the amplitudes $\alpha(z)$ are given from the weights of two edges out from each node on the path determined by a given bit sequence $z$; the amplitudes are restricted, and dependent on the support of the Boolean function of the $\mathrm{WBDD}_{f}$ since the $\mathrm{WBDD}_{f}$ assumes by its contraction rules that equivalent subgraphs are shared, and that deleted nodes have the same edge weights. In general, the probabilities independent of the support need another computational resource, for instance, a quantum random access memory (QRAM), that stores quantum accessible classical bits. In the implicit model of QRAM in Ref. Matteo et al. (2019), locations in memory are addressed by $n$-bit strings, and are queried by inputting the associated state to a circuit. In the BDD case, using $z_{i}$-edges on a path as an access key to QRAM, we can represent a Boolean function, if the function can be decomposed into the sequence of QRAM values accessed by sequential keys of sequential paths $P_{1}\subset P_{2}\subset\cdots$. (See also Appendix A.)
Following Ref. Breitbar P et al. (1995), we give an example of an $O(\mathrm{poly}(n))$-sized $\mathrm{FBDD}_{f}$ that is a classical description of a quantum state with exponentially large nonzero coefficients. First, let us define the following symmetric Boolean function.
Definition 6.
Let $X=(x_{1},\ldots,x_{n})$, and $S^{i}_{n}(X)$ be $n$ variables, and the symmetric Boolean function that is equal to one if and only if exactly $0\leq i\leq n$ of its variables are equal to one. The function $S^{i}_{n}(X)$ is called the elementary symmetric function.
Second, we define the following function (see Fig. 5),
$f^{1}(X,Y,Z)$ $=$ $\sum_{i=1}^{n}x_{i}(S^{i-1}_{2n}(Y,Z)+S^{n+i}_{2n}(Y,Z)),$ (9)
$f^{2}(X,Y,Z)$ $=$ $f^{1}(Y,Z,X),$ (10)
$f^{3}(X,Y,Z)$ $=$ $f^{1}(Z,X,Y).$ (11)
Then, we define the following Boolean function $h$ (see Fig. 5),
$h(X,Y,Z)=\bar{v}\bar{w}f^{1}(X,Y,Z)+\bar{v}wf^{2}(X,Y,Z)+v\bar{w}f^{3}(X,Y,Z).$ (12)
The following result (Theorem 7 of Ref. Breitbar P et al. (1995)) is known. For every $n\geq 4$, the $\mathrm{FBDD}_{h}$ is computable with no more than $33(2n^{2}+2n)+$ 3 nodes; however, every OBDD computing $h$ needs at least $2^{n/3}$ nodes. Thus, if the number of $z$ such as $h(z)=1$ is $O(poly(n))$, then the size of an OBDD computing $h$ is $O(poly(n))$, which contradicts Theorem 7 of Ref. Breitbar P et al. (1995). Therefore, the $\mathrm{FBDD}_{h}$ with weighted edges is a case of an $n$-qubit quantum state with $d=O(2^{n})$ nonzero coefficients.
Figure 5: (a) A $6$-variable example of subBDD for $f^{1}(X,Y,Z)=x_{1}(S_{4}^{0}(Y,Z)+S_{4}^{3}(Y,Z))+x_{2}(S_{4}^{1}(Y,Z)+S_{4}^{4}(Y,Z))$, where $X=(x_{1},x_{2}),Y=(y_{1},y_{2}),Z=(z_{1},z_{2})$. The function $S_{n}^{i}(x_{1},\ldots,x_{n})$ is called the elementary symmetric function that is one if and only if exactly $0\leq i\leq n$ of its variables are equal to one. (b) A $\mathrm{FBDD}_{h}$ for $h(X,Y,Z)=\bar{v}\bar{w}f^{1}(X,Y,Z)+\bar{v}wf^{2}(X,Y,Z)+v\bar{w}f^{3}(X,Y,Z)$.
III.2 Quantum algorithm for QSP from weighted FBDD
Now, we show that we can construct an efficient quantum circuit generating the quantum state described by a given $\mathrm{WFBDD}_{f}$ (see an example in Fig. 6).
Algorithm 1 StateGeneration
1:A $\mathrm{WFBDD}_{f}=(V,E)$ in Def. 4, where let $i(u)$, $e_{b}(u)$, $h_{b}(u)$, and $w_{b}(u)$ denote the index of the variable assigned to an internal node $u$, the $b$-edge out from $u$, the head node of $e_{b}(u)$, and the weight assigned to $e_{b}(u)$, respectively.
2:A quantum state $\sum_{z\in\{0,1\}^{n}}\alpha(z)\ket{z}$ described by the $\mathrm{WFBDD}_{f}$ in Eq. (1).
3:[Initialization]
4:Prepare a set $Q=\{r\}$, and an empty stack $S$.
5:Assign qubits set to $\ket{+}$ to the variables $x_{1},\ldots,x_{n}$.
6:Assign a qubit set to |$\ket{1}$ to the root $r$.
7:Assign qubits set to |$\ket{0}$ to the internal nodes except the root $r$.
8:[Main loop]
9:while the set $Q$ is not empty do
10:     Remove a node $u$ from $Q$.
11:     Push the node $u$ to the stack $S$.
12:     Apply the control-$H$ gate controlled by the qubit $u$ to the qubit $x_{i(u)}$.
13:     Apply the control-$U(u)$ gate controlled by the qubit $u$ to the qubit $x_{i(u)}$, where $U(u)=\frac{1}{\sqrt{|w_{0}(u)|^{2}+|w_{1}(u)|^{2}}}\begin{pmatrix}w_{0}(u)&-w^{\ast}_{1}(u)\\ w_{1}(u)&w^{\ast}_{0}(u)\\ \end{pmatrix}$.
14:     for each $b$-edge $e_{b}(u)$ out from the node $u$, and its head node $h_{b}(u)$ do
15:         Remove the $b$-edge $e_{b}(u)$ from $E$.
16:         if the head node $h_{b}(u)$ has no incoming edge, and is not the terminal node then
17:              Add the head node $h_{b}(u)$ to the set $Q$.
18:         end if
19:         if $h_{b}(u)\neq t_{0},t_{1}$ then
20:              Apply the $X$ gate to the qubit $x_{i}$ when $b=0$.
21:              Apply the Toffoli gate controlled by the qubits $u$, and $x_{i(u)}$ to the qubit $h_{b}(u)$.
22:              Apply the $X$ gate on the register assigned to $x_{i}$ when $b=0$.
23:         end if
24:     end for
25:end while
26:[Undo]
27:while the stack $S$ is not empty do
28:     Pop a node $u$ from the stack $S$.
29:     Do Lines 14, 1923.
30:end while
Theorem 1.
For a given $\mathrm{WFBDD}_{f}=(V,E)$ encoding an $n$-bit Boolean function $f$, Algorithm 1 gives the quantum state described by the $\mathrm{WFBDD}_{f}$, with $O(|V|)$ uses of single- or two-qubit gates, and $|V|-2$ ancillary qubits.
Figure 6: An example of Algorithm 1. (a) An example of $\textrm{WFBDD}_{f}$. In Algorithm 1, all the internal nodes are pushed to a stack $S$ in a topological ordering such that for every directed edge $(u,v)$ out from $u$ to $v$, $u$ comes before $v$, i.e., $r\prec u_{1}\prec\cdots\prec u_{5}$. Start from the root $r$, and push $r$ to $S$. Then, remove the 00-edge $(r,u_{1})$, and the $1$-edge $(r,u_{2})$, at which, if the indegree of $u_{1(2)}$ is zero, $u_{1(2)}$ is added to a set $Q$. The added nodes are pushed to $S$ in the following rounds, and ordered. Repeating this process, we push all the internal nodes to $S$, and ordered. (b) The quantum circuit generated from the $\textrm{WFBDD}_{f}$. Initially, assign qubits to the variables $x_{1},\ldots,x_{4}$, and the internal nodes $r,u_{1},\ldots,u_{5}$. Each qubit $x_{i}$ is set to $\ket{+}$, the qubit $r$ is set to |$\ket{1}$, and each qubit $u_{j}$ is set to |$\ket{0}$. Then, the control-$H$ gate, and the control-$U(r)$ gate (see Eq. (15)) controlled by the qubit $r$ are applied to the qubit $x_{1}$. Furthermore, the Toffoli gate controlled by the qubits $r$, and $x_{1}$ is applied to each qubit $u_{1(2)}$ as shown in the figure. Preparing the unitaries for rest of the internal nodes in the topological order, we derive the quantum circuit in the figure.
Proof.
First of all, we show that Algorithm 1 gives a stack $S$ of all internal nodes of $\textrm{WFBDD}_{f}$ with a linear total ordering such that for every directed edge $(u,v)$ out from an internal node $u$ to another internal node $v$, $u$ comes before $v$ in the ordering, i.e., a topological sort. In Algorithm 1, a set $Q=\{r\}$ consisting of the root $r$ is prepared (Line 4), $r$ is removed from $Q$ (Line 10), and pushed to $S$ (Line 11). Since $r$ is indegree 00, $S$ follows the ordering. Next, for each $b$-edge $e_{b}(r)$ out from $r$ to its head node $h_{b}(r)$, we remove $e_{b}(r)$ from $E$, and check the indegree of $h_{b}(r)$, and whether $h_{b}(r)$ is the terminal nodes $t_{0},t_{1}$ or not (Lines 14-16). If the indegree is zero, and $h_{b}(u)\neq t_{0},t_{1}$, then $h_{b}(r)$ is added to $Q$ (Line 17). The added node $h_{b}(r)$ is removed from $Q$, and pushed to $S$ in some round of the while loop. The head node $h_{b}(r)$ follows the ordering because the indegree zero means that all the directed edges whose head node is $h_{b}(r)$ have been removed from $E$, and the tail nodes of the directed edges have been pushed to $S$. Repeating this process, we order all the nodes. If there is an unordered internal node $u$, then the indegree of $u$ must be nonzero, which means another unordered internal node. Repeating this discussion leads that the root $r$ is not unordered, which contradicts that $r$ is first pushed to $S$. Let $\prec$ denote the total ordering.
We show that Algorithm 1 gives the following state by induction:
|$\ket{\psi_{u}}=\sum_{z\in\{0,1\}^{n}}\alpha_{u}(z)\ket{z}\ket{g_{u}(z)},$ (13)
where $\alpha_{u}(z)$, and $g_{u}(z)$ are defined as follows. Define $V_{u}=\{v\in V\ |\ v\preceq u\}$ as a set of an internal node $u$, and all the preceding nodes to $u$, and $P(z)=(V(z),E(z))$ as the path derived by starting from the root $r$, and selecting the $z_{i}$-edge in each internal node until reaching the terminal node $t_{1}$. Then, $\alpha_{u}(z)$ is defined as
$\alpha_{u}(z)=\frac{1}{\sqrt{2^{n-|V_{u}\cap V(z)|}}}\prod_{v\in V_{u}\cap V(z)}\frac{w_{z_{i(v)}}(v)}{\sqrt{|w_{0}(v)|^{2}+|w_{1}(v)|^{2}}}.$ (14)
We define $g_{u}(z)$ as the $(|V|-2)$-bit sequence such that each bit is assigned to each node $v\in V$ except the terminal nodes $t_{0}$, and $t_{1}$, and each bit value assigned to the node $v$ is $1$ if $v\in(V_{u}\cup\{h_{0}(u),h_{1}(u)\})\cap V(z)$, where $h_{b}(u)$ is the head node of the $b$-edge $e_{b}(u)$ of $u$ without the terminal nodes $t_{0},t_{1}$. When $u=r$, the control-$H$ gate, and the control-$U(r)$ gate controlled by the qubit $r$ set to |$\ket{1}$ are applied to the qubit $x_{i(r)}$ set to $\ket{+}$ (Lines 5, 6, 12, and 13), where
$U(r)=\frac{1}{\sqrt{|w_{0}(r)|^{2}+|w_{1}(r)|^{2}}}\begin{pmatrix}w_{0}(r)&-w^{\ast}_{1}(r)\\ w_{1}(r)&w^{\ast}_{0}(r)\\ \end{pmatrix},$ (15)
, and $w_{b}(r)$ is the weight assigned to $e_{b}(u)$. Thus, Eq. (14) is satisfied for the root $r$,
$\alpha_{r}(z)=\frac{1}{\sqrt{2^{n-1}}}\frac{w_{z_{i(r)}}(r)}{\sqrt{|w_{0}(r)|^{2}+|w_{1}(r)|^{2}}}.$ (16)
Furthermore, for the head node $h_{0}(r)\neq t_{0},t_{1}$, the qubit $x_{i(r)}$ is flipped, and the Toffoli gate controlled by the qubits $r$, and $x_{i(r)}$ is applied to the qubit $h_{0}(r)$. Then, the qubit $x_{i(r)}$ is flipped again (Lines 19-22). For the head node $h_{1}(r)\neq t_{0},t_{1}$, the same operation is also performed except the two flipping operations. Remembering that the qubit $r$ is set to |$\ket{1}$ (Line 6), we derive that $g_{r}(z)$ satisfies its definition.
Assume that $\alpha_{u}(z)$, and $g_{u}(z)$ satisfy their definitions for an internal node $u$, and any internal node preceding to the node, and consider the internal node $v$ following the node $u$ in the ordering. For an arbitrary path $P(z)=(V(z),E(z))$ from the root $r$, through $v$, and to the terminal node $t_{1}$, let $e_{b}(u)=(u,v)\in E(z)$, and $e_{b^{\prime}}(v)=(v,w)\in E(z)$ denote directed edges into, and out from the node $v$, respectively. Since $g_{u}(z)$ satisfies its definition and thus the qubit $v$ is set to |$\ket{1}$, Applying both of the controlled-$H$ gate and the controlled-$U(v)$ gate, controlled by the qubit $v$, to qubit $x_{i(v)}$ set to $\ket{+}$ (Lines 5, 12, and 13), we obtain that
$\alpha_{u}(z)\frac{\sqrt{2}w_{z_{i(v)}}(v)}{\sqrt{|w_{0}(v)|^{2}+|w_{1}(v)|^{2}}}=\frac{1}{\sqrt{2^{n-|V_{v}\cap V(z)|}}}\prod_{v^{\prime}\in V_{v}\cap V(z)}\frac{w_{z_{i(v^{\prime})}}(v^{\prime})}{\sqrt{|w_{0}(v^{\prime})|^{2}+|w_{1}(v^{\prime})|^{2}}}=\alpha_{v}(z).$ (17)
Furthermore, when $b^{\prime}=0$, the qubit $x_{i(v)}$ is flipped, and the Toffoli gate controlled by the qubits $v$, and $x_{i(v)}$ is applied to the qubit $w$, which is the head node $h_{b^{\prime}}(v)$ of the $b^{\prime}$-edge $e_{b^{\prime}}(v)$. Then, the qubit $x_{i(v)}$ is flipped again (Lines 19-22). When $b^{\prime}=1$, the same operation is also performed except the two flipping operations. Thus, $g_{v}(z)$ also satisfies its definition.
The gate complexity is derived trivially from Algorithm 1. Note that the number of uses of Toffoli gates is at most $|E|=2|V|$ because a binary decision diagram has two edges per node. ∎
IV Quantum state preparation for uniform superpositions
In the previous section, we showed that an arbitrary $\mathrm{WFBDD}_{f}$ is a classical description of an $n$-qubit quantum state whose support is {|$\{\ket{x}\ |\ x\in\{0,1\}^{n},f(x)=1\}$, and proposed a classical algorithm to construct a quantum circuit to prepare the quantum state. In this section, we explore to prepare a uniform superposition from a Boolean function $f:x\in\{0,1\}^{n}\mapsto f(x)\in\{0,1\}$. We define a quantum state encoding the Boolean function $f$ as
|$\ket{\phi_{f}}=\frac{1}{\sqrt{|f|}}\sum_{x\in\{0,1\}^{n}}f(x)\ket{x},$ (18)
where $|f|$ is the number of $x\in\{0,1\}^{n}$ such that $f(z)=1$, i.e., the amplitude encoding of the Boolean function. Then, out task is to assign appropriate weights to convert a given $\textrm{FBDD}_{f}$ to the weighted FBDD which describes a quantum state encoding the Boolean function $f$ (see an example in Fig. 7).
Theorem 2.
For a given $\textrm{FBDD}_{f}=(V,E)$ encoding an $n$-bit Boolean function $f$, Algorithm 2 with at most $26|V|-$ 2 queries to the nodes in $V$ converts the $\textrm{FBDD}_{f}$ to one $\textrm{WFBDD}_{f}$ which describes a quantum state encoding $f$.
Algorithm 2 UniformWeightDistribution
1:A $\textrm{FBDD}_{f}=(V,E)$ encoding a Boolean function $f(x_{1},\ldots,x_{n})$ in Def. 2, where let $i(u)$, $e_{b}(u)$, and $h_{b}(u)$ denote the index of the variable assigned to an internal node $u$, the $b$-edge out from $u$, and the head node of $e_{b}(u)$, respectively.
2:A $\textrm{WFBDD}_{f}$ which describes a quantum state encoding $f$, where let $w_{b}(u)$ denote the weight assigned to $e_{b}(u)$.
3:[Initialization]
4:$X_{u}:=\emptyset$ for each internal node $u$.
5:$X_{t_{0}}:=\{x_{1},\ldots,x_{n}\}$, and $X_{t_{1}}:=\{x_{1},\ldots,x_{n}\}$, where $t_{0}$, and $t_{1}$ are the terminal nodes.
6:$w_{0}(t_{0}):=0$, $w_{1}(t_{0}):=0$, $w_{0}(t_{1}):=0$, and $w_{1}(t_{1}):=1$ conveniently.
7:Enqueue $t_{1}$ to a queue $Q$.
8:[Main loop]
9:while $Q$ is not empty do
10:     Dequeue a node $v$ from $Q$.
11:     for each tail node $u$ of all the directed edges into $v$ do
12:         if $X_{h_{0}(u)}\neq\emptyset$, and $X_{h_{1}(u)}\neq\emptyset$ then
13:              Enqueue the node $u$ to $Q$.
14:              $X=X_{h_{0}(u)}\cap X_{h_{1}(u)}$.
15:              for each $h_{b}(u)$ do
16:                  Assign $\sqrt{w}$ to the weight $w_{b}(u)$ of $e_{b}(u)$, where $w=(w^{2}_{0}(h_{b}(u))+w^{2}_{1}(h_{b}(u)))2^{|X_{h_{b}(u)}\setminus X|}$.
17:              end for
18:              $X_{u}=X\setminus\{x_{i(u)}\}$.
19:         end if
20:     end for
21:end while
Figure 7: An example of Algorithm 2. (a) Initialization. For the terminal nodes $t_{0}$, and $t_{1}$, define the set $X_{t_{0}}=X_{t_{1}}:=\{x_{1},x_{2},x_{3},x_{4}\}$ of unassigned variables, and assign 00 to all the directed edges into $t_{0}$. (b) Starting from $t_{1}$, take nodes $u$ satisfying the followings. One is that $u$ is a tail node of the $b$-edge of $t_{1}$. The other is that any unassigned variables $X_{h_{b}(u)}$ is not empty, where $h_{b}(u)$ is the head node of the $b$-edge of $u$. In this round, $v_{1}$, and $v_{2}$ satisfy the conditions. For $v_{1}$, since $X=X_{t_{0}}\cap X_{t_{1}}=\{x_{1},x_{2},x_{3},x_{4}\}$, and $X_{t_{1}}\setminus X=\emptyset$, there is no deleted nodes between $v_{1}$, and $t_{1}$, and thus the weight $w_{0}(v_{1})$ is assigned $1$. Since $x_{3}$ is assigned to $v_{1}$, $X_{v_{1}}$ is $X\setminus\{x_{3}\}=\{x_{1},x_{2},x_{4}\}$. The weight $w_{1}(v_{2})$ is assigned $1$ in the same way. (c) Similarly, take nodes $v_{3},v_{4}$ satisfying the two conditions in (b). For $v_{3}$, and the $1$-edge $e_{1}(v_{3})$ of $v_{3}$, since $X=X_{v_{1}}\cap X_{t_{1}}=\{x_{1},x_{2},x_{4}\}$, and $X_{t_{1}}\setminus X=\{x_{3}\}$, there is one deleted node assigned $x_{3}$. Thus, the weight $w_{1}(v_{3})$ is assigned $\sqrt{w}=\sqrt{2}$, where $w$ is calculated by $w=(w^{2}_{0}(v_{1})+w^{2}_{1}(t_{1}))2^{|X_{t_{1}}\setminus X|}=2$. The weight $w_{0}(v_{3})$ is assigned $1$ in the same way. (d), (e) Repeating the above process, we assign the unassigned variables and the weights to all the nodes and all the directed edges in self-consistency, respectively.
Proof.
First, we show that Algorithm 2 gives a linear total ordering with which all the nodes follow. In Algorithm 2, whether the set $X_{u}$ is empty serves as the criterion for determining the order of the node $u$. Thus, for each internal node $u$, $X_{u}$ is initialized to be an empty set, and for the terminal nodes $t_{0}$, and $t_{1}$, $X_{t_{0}}=X_{t_{1}}:=\{x_{1},\ldots,x_{n}\}$ (Lines 4, and 5). The terminal node $t_{1}$ is enqueued to a queue $Q$ (Line 7). For each tail node $u$ of all the directed edges into $t_{1}$, check whether both of $X_{h_{0}(u)}$, and $X_{h_{1}(u)}$ are non-empty or not (Lines 11, and 12), where $h_{b}(u)$ is the head node of the directed edge. If both of $X_{h_{0}(u)}$, and $X_{h_{1}(u)}$ are non-empty, $u$ is enqueued to $Q$, and $X_{u}$ is updated to be $X_{u}=(X_{h_{0}(u)}\cap X_{h_{1}(u)})\setminus\{x_{i(u)}\}$, where $i(u)$ is the index of the variable assigned to $u$ (Lines 12-14, and 18). Repeating this process, we order all the nodes, since if there exists a non-ordered node $u$, then at least either of $X_{h_{0}(u)}$ or $X_{h_{1}(u)}$ is empty which derives the contradiction that either of $X_{t_{0}}$ or $X_{t_{1}}$ is empty finally. Let $\prec$ denote the total ordering.
Second, we show that Algorithm 2 assigns self-consistent subset of variables $X_{u}$ for each node $u$. FBDDs by definition have at most one occurrence of each variable for each path, and then we can consider that each edge on the path contains hidden label variables. In the case of OBDDs, hidden label variables at each edge $(u,v)$ are given by checking indices $i(u)$, and $i(v)$ because the order of occurrences of variables of each path is consistent with some linear order. However, this check does not work for the FBDDs, because the FBDDs have a different order of occurrences of variables on each path. Thus, we have to assign variables to each edge self-consistently.
By induction, we show that for each node $u$, $X_{u}$ is a set of variables which are unassigned on the path from the root $r$ to $u$ yet. For the terminal node $t_{b}$, $X_{t_{b}}$ is initialized to be $X_{t_{b}}=\{x_{1},\ldots,x_{n}\}$ (Line 5), and it satisfies the claim clearly. Then, assume that all the nodes $u^{\prime}\preceq u$ satisfies the claim, and take a node $v$ such that for the head node $h_{b}(v)$ of the $b$-edge $e_{b}(v)$ of $v$, $h_{b}(v)\preceq u$. From the assumption, both of $X_{h_{0}(v)}$, and $X_{h_{1}(v)}$ are not empty, and are the sets of unassigned variables on paths from $r$ to $h_{0}(v)$ or $h_{1}(v)$ respectively. Thus, assigning variables $X_{h_{b}(v)}\setminus(X_{h_{0}(v)}\cap X_{h_{1}(v)})$ to hidden nodes between $v$, and $h_{b}(v)$, we derive that it is self-consistent since the common variables of $X_{h_{0}(v)}$, and $X_{h_{1}(v)}$ are postponed to be assigned to the nodes on paths from $r$ to $v$.
Finally, consider a directed edge $(u,v)\in E$, and assume that the head node $h_{0}(u)=v$ of the edge without loss of generality. Remembering that $w_{0}(u)$ is calculated by
$w_{0}(u)=(w^{2}_{0}(v)+w^{2}_{1}(v))2^{|X_{v}\setminus X|},$ (19)
where $X=(X_{v}\cap X_{h_{1}(u)})$ (Line 16), we derive that
$\frac{w_{0}(u)}{\sqrt{w^{2}_{0}(u)+w^{2}_{1}(u)}}\frac{1}{\sqrt{w^{2}_{0}(v)+w^{2}_{1}(v)}}=\sqrt{\frac{2^{|X_{v}\setminus X|}}{w^{2}_{0}(u)+w^{2}_{1}(u)}}.$ (20)
Thus, for a given $z\in\{0,1\}^{n}$ such that $f(z)=1$, the corresponding coefficient $\alpha(z)$ is calculated by
$\alpha(z)$ $=$ $\frac{1}{\sqrt{2^{n-|V_{z}|+1}}}\prod_{u\in V_{z}\setminus\{t_{0},t_{1}\}}\frac{w_{z_{i(u)}}(u)}{\sqrt{|w_{0}(u)|^{2}+|w_{1}(u)|^{2}}}$ (21)
$=$ $\frac{1}{\sqrt{2^{n-|V_{z}|+1}}}\sqrt{\frac{2^{n-|V_{z}|+1}}{|w_{0}(r)|^{2}+|w_{1}(r)|^{2}}}=\frac{1}{\sqrt{|w_{0}(r)|^{2}+|w_{1}(r)|^{2}}}.$ (22)
In Algorithm 2, all the nodes except $t_{0}$ are enqueued to (and dequeued from) $Q$, and for each node $u$, all the in-edges, and out-edges are processed. Thus, the query complexity to the nodes in $\mathrm{FBDD}_{f}$ is $22(|V|-1)+2|E|\leq 6|V|-$ 2. ∎
V Controlled Phase operation based on FBDD
In this section, we show another encoding of a Boolean function to a quantum state. For a given ${\rm FBDD}_{f}$, we show that Algorithm 3 constructs a quantum circuit realizing a unitary such that $U_{f}\ket{x}=e^{i\theta f(x)}\ket{x}$. Using such a quantum circuit, we can trivially prepare a $n$-qubit quantum state
|$\ket{\psi_{\theta}}:=\frac{1}{\sqrt{2^{n}}}\sum_{x\in\{0,1\}}e^{i\theta f(x)}\ket{x},$ (23)
from a uniform superposition $(H\ket{0})^{\otimes n}$.
Algorithm 3 is derived by modifying Algorithm 1 slightly. In Algorithm 1, notice that the edge-label sequence $z=z_{1}\cdots z_{n}$ of a path from the root $r$ to the terminal node $t_{1}$ represents the |$\ket{z}$ in the support {|$\{\ket{x}\ |\ x\in\{0,1\}^{n},f(x)=1\}$, and the nodes on the path are recorded in the ancillary qubits. Then, to construct the quantum circuit realizing a unitary such that $U_{f}\ket{x}=e^{i\theta f(x)}\ket{x}$, first, let $n$ qubits assigned to the variables $x_{1},\ldots,x_{n}$ just be the input quantum registers (or replace Line 5 of Algorithm 1 with Line 5 of Algorithm 3), and assign an ancillary qubit to the terminal node $t_{1}$ on which the result $f(x)$ are output (Line 8 of Algorithm 3). Second, remove the controlled-$H$ gates, and controlled-$U(u)$ gates (or remove Lines 12, and 13 of Algorithm 1), and remove the terminal node $t_{1}$ from the condition in Line 19 (or see Line 18 of Algorithm 3) that result in adding Toffoli gates to output $f(x)$ to the last ancillary qubit $t_{1}$. Finally, apply a phase gate to the ancillary qubit $t_{1}$ (Line 25), and undo all the Toffoli gates. (See also Fig. 8 as an example of Algorithm 3.) We give a corollary as a summary of the above discussion.
Corollary 1.
For a given ${\rm FBDD}_{f}=(V,E)$ encoding a $n$-bit Boolean function $f$, Algorithm 3 gives a quantum circuit $U_{f}$ consisting of a phase gate, and at most $2|V|$ Toffoli gates with $|V|-1$ ancillary qubits such that for $x\in\{0,1\}^{n}$,
$U_{f}:\ket{x}\otimes\ket{0}^{\otimes|V|-1}\mapsto e^{i\theta f(x)}\ket{x}\otimes\ket{0}^{\otimes|V|-1}.$ (24)
Algorithm 3 ControlledPhase
1:A $\mathrm{FBDD}_{f}=(V,E)$ in Def. 2, where let $i(u)$, $e_{b}(u)$, and $h_{b}(u)$ denote the index of the variable assigned to an internal node $u$, the $b$-edge out from $u$, and the head node of $e_{b}(u)$, respectively.
2:A quantum circuit realizing a unitary such that $U_{f}\ket{x}=e^{i\theta f(x)}\ket{x}$.
3:[Initialization]
4:Prepare a set $Q=\{r\}$, and an empty stack $S$.
5:Assign qubits for an input state to the variables $x_{1},\ldots,x_{n}$.
6:Assign a qubit set to |$\ket{1}$ to the root $r$.
7:Assign qubits set to |$\ket{0}$ to the internal nodes except the root $r$.
8:Assign a qubit set to |$\ket{0}$ to the terminal node $t_{1}$.
9:[Main loop]
10:while the set $Q$ is not empty do
11:     Remove a node $u$ from $Q$.
12:     Push the node $u$ to the stack $S$.
13:     for each $b$-edge $e_{b}(u)$ out from the node $u$, and its head node $h_{b}(u)$ do
14:         Remove the $b$-edge $e_{b}(u)$ from $E$.
15:         if the head node $h_{b}(u)$ has no incoming edge, and is not the terminal node then
16:              Add the head node $h_{b}(u)$ to the set $Q$.
17:         end if
18:         if $h_{b}(u)\neq t_{0}$ then
19:              Apply the $X$ gate to the qubit $x_{i}$ when $b=0$.
20:              Apply the Toffoli gate to the qubit $h_{b}(u)$ controlled by the qubits $u$, and $x_{i(u)}$.
21:              Apply the $X$ gate on the register assigned to $x_{i}$ when $b=0$.
22:         end if
23:     end for
24:end while
25:Apply a phase gate to the qubit $t_{1}$.
26:[Undo]
27:while the stack $S$ is not empty do
28:     Pop a node $u$ from the stack $S$.
29:     Do Lines 1323.
30:end while
Figure 8: An example of Algorithm 3. (a) An example of $\textrm{FBDD}_{f}$. In Algorithm 3, all the internal nodes are pushed to a stack $S$ in a topological ordering such that for every directed edge $(u,v)$ out from $u$ to $v$, $u$ comes before $v$, i.e., $u\prec u_{1}\prec\cdots\prec u_{5}$. Start from the root $r$, and push $r$ to $S$. Then, remove the 00-edge $(r,u_{1})$, and the $1$-edge $(r,u_{2})$, at which, if the indegree of $u_{1(2)}$ is zero, $u_{1(2)}$ is added to a set $Q$. The added nodes are pushed to $S$ in the following rounds, and ordered. Repeating this process, we order all the internal nodes, and push them to $S$. (b) The quantum circuit generated from the $\textrm{FBDD}_{f}$. Initially, assign qubits to the variables $x_{1},\ldots,x_{4}$, the internal nodes $r,u_{1},\ldots,u_{5}$, and the terminal node $t_{1}$. Each qubit $x_{i}$ is for the input qubit, the qubit $r$ is set to |$\ket{1}$, and each qubit $u_{j}$ is set to |$\ket{0}$. Then, the Toffoli gate controlled by the qubits $r$, and $x_{1}$ is applied to each qubit $u_{1(2)}$ as shown in the figure. Applying the Toffoli gates for rest of the internal nodes in the topological ordering, we derive the quantum circuit in the figure.
VI Application to Black-Box QSP
In this section, we give an $n$-qubit quantum state that cannot be efficiently prepared by the black-box QSP Bausch (2022) but can be generated efficiently by our proposed QSP. For a given $\delta\in(0,1)$, we consider an $n$-qubit quantum state
|$\ket{\psi_{\delta}}=\frac{1}{||\alpha||_{2}}\sum_{x\in\{0,1\}^{n}}\alpha_{x}\ket{x},$ (25)
where $\alpha_{x}=\delta^{w_{H}(x)}/\binom{n}{w_{H}(x)}$, and $w_{H}(x)$ is the Hamming weight of $x$. We show that the number of the amplitude amplifications to generate the quantum state |$\ket{\psi_{\delta}}$ cannot be reduced exponentially by the black-box QSP, and we give a weighted FBDD describing |$\ket{\psi_{\delta}}$.
The black-box QSP in Ref. Bausch (2022) transforms an initial state through an intermediate state to the target state by performing the nested amplitude amplifications. Furthermore, for an initial state, and a target state
|$\ket{s}$ ::=:= $\frac{1}{\lVert\bar{A}\rVert_{2}}\sum_{j=0}^{g-1}\bar{A}_{j}\ket{j},$ (26)
|$\ket{t}$ ::=:= $\frac{1}{\lVert\alpha\rVert_{2}}\sum_{i=0}^{N-1}\alpha_{i}\ket{i},$ (27)
where
$\bar{A}_{j}$ ::=:= $\sum_{i=0}^{N-1}b_{ij}\sqrt{w_{j}},$ (28)
$\alpha_{i}$ ::=:= $\sum_{j=0}^{g-1}b_{ij}w_{j},$ (29)
and $B=(b_{ij})$ is some Boolean matrix, the number of the nested amplitude amplifications is shown to be evaluated as $O\left(\sqrt{N}\lVert\alpha\rVert_{1}/\lVert\bar{A}\rVert_{2}\right)$ Bausch (2022). We write that $N=2^{n}$, $g=n+1$, and
$b_{ij}$ $=$ $\left\{\begin{array}[]{ll}1&w_{H}(i)=j,\\ 0&\text{the others},\end{array}\right.$ (32)
$w_{j}$ $=$ $\frac{\delta^{j}}{\binom{n}{j}},$ (33)
we derive that
$\bar{A}_{j}$ $=$ $\left(\binom{n}{j}\delta^{j}\right)^{\frac{1}{2}},$ (34)
$\alpha_{i}$ $=$ $\frac{\delta^{w_{H}(i)}}{\binom{n}{w_{H}(i)}},$ (35)
and then
$\sqrt{N}\frac{\lVert\alpha\rVert_{1}}{\lVert\bar{A}\rVert_{2}}=\sqrt{\frac{1-\delta^{n+1}}{1-\delta}}\left(\frac{2}{1+\delta}\right)^{\frac{n}{2}}.$ (36)
Thus, the number of nested amplitude amplifications to generate the quantum state |$\ket{\psi_{\delta}}$ cannot be reduced exponentially.
Figure 9: A weighted OBDD describing a $4$-qubit quantum state in Eq. (25). The weighted OBDD consists of $4$ layers of which the $i$-th layer has $i$ nodes labeled with a variable $x_{i}$. For $i=1,2,3$, and $j\in[i]$, the $j$-th node in the $i$-th layer connects to the $j$-th node in the $(i+1)$-th layer with the $1$-edge, and connects to the $(j+1)$-th node in the $(i+1)$-th layer with the 00-edge. Any node in the $4$-th layer connects to the $1$-terminal node with both of the 00-edge, and the $1$-edge. Weights are assigned as follows. For example, for an edge $e_{1}$ going into the terminal node $t_{1}$, notice that any path from the root $r$ to $t_{1}$ through the edge $e_{1}$ pass through three $1$-edges. Thus, remembering that the coefficient $\alpha_{x}$ is determined by the Hamming weight $w_{H}(x)$, we can assign $\alpha_{x}=\delta^{3}/4$ on the edge $e_{1}$. A weight of the other unweighted edge, e.g., $e_{2}$, is the root of the sum $\left(\delta^{4}/36+\delta^{6}/16\right)^{1/2}$ by two squares of the weights of the edges outgoing from the node $v$ into which $e_{2}$ goes.
On the other hand, we show that the $n$-qubit quantum state |$\ket{\psi_{\delta}}$ can be written in the form of a weighted OBDD with $n(n+1)/2$ nodes. Then, from Theorem 1, Algorithm 1 gives the quantum state |$\ket{\psi_{\delta}}$ with $O(n(n+1))$ uses of single- or two-qubit gates, and $n(n+1)/2-2$ ancillary qubits.
The weighted OBDD consists of $n$ layers of which the $i$-th layer has $i$ nodes labeled with a variable $x_{i}$. (See Fig. 9 in the case of $4$ qubits.) For $i\in[n-1]$, and $j\in[i]$, the $j$-th node in the $i$-th layer connects to the $j$-th node in the $(i+1)$-th layer with the $1$-edge, and connects to the $(j+1)$-th node in the $(i+1)$-th layer with the 00-edge. Any node in the $n$-th layer connects to the $1$-terminal node with both of the 00-edge, and the $1$-edge.
Weights are assigned as follows. First, for an edge $e$ going into the terminal node $t_{1}$, notice that any path from the root $r$ to $t_{1}$ through the edge $e$ pass through the same number of the $1$-edges. Thus, remembering that the coefficient $\alpha_{x}$ is determined by the Hamming weight $w_{H}(x)$, we can assign each $\alpha_{x}$ on each edge going into the terminal node $t_{1}$ based on the Hamming weight $w_{H}(x)$ of $x$, and the number of the passed $1$-edges. The weight of the other unweighted edge is the root of the sum by two squares of the edges outgoing from the node into which the unweighted edge goes.
VII Application to the block encoding
Our techniques for QSP can be applied to the block encoding, a technique of encoding a matrix as a block of a unitary, defined in the following (Definition 47 of Ref. Gilyén et al. (2018)).
Definition 7.
Suppose that $A$ is an $s$-qubit operator, $\alpha,\epsilon\in\mathbb{R}_{+}$, and $a\in\mathbb{N}$, then we say that the $(s+a)$-qubit unitary $U$ is an $(\alpha,a,\epsilon)$-block encoding of $A$, if
$||A-\alpha(\bra{0}^{\otimes a}\otimes I)U(\ket{0}^{\otimes a}\otimes I)||\leq\epsilon,$ (37)
where |||||$\cdot$||| | ⋅ | | is the operator norm. Note that since $||U||=1$, we necessarily have $||A||\leq\alpha+\epsilon$.
References Gilyén et al. (2018, 2019) give the instructions of constructing block encodings of density operators, POVM operators, Gram matrices, and sparse-access matrices, respectively (Lemma 45 - 48 of Ref. Gilyén et al. (2018)). Therefore, when the unitary operators or the oracles assumed in each lemma of Ref. Gilyén et al. (2018) are given through weighted FBDDs, we can construct the block encodings.
For example, we consider a $\mathrm{WFBDD}_{f}$ representing an $n$-qubit quantum state |$\ket{\psi_{f}}=\sum_{z}c_{z}f(z)\ket{z}$. From Algorithm 1, we can implement a quantum circuit $U_{f}$ such that
$U_{f}:\ket{0}^{\otimes n}\otimes\ket{0}^{\otimes|V|-2}\mapsto\sum_{z\in\{0,1\}^{n}}c_{z}f(z)\ket{z}\otimes\ket{0}^{\otimes|V|-2},$ (38)
where $|V|$ is the number of nodes. Then, $(U_{f}^{\dagger}\otimes I_{n})\textsc{SWAP}(U_{f}\otimes I_{n})$ is the $(1,n+|V|-2,0)$-block encoding of |$\ket{\psi_{f}}\bra{\psi_{f}}$ |, where SWAP is the operation swapping the first $n$ qubits, and the last $n$ qubits. As shown in Lemma 45 of Ref. Gilyén et al. (2019), we have that
$\bra{0}^{\otimes n+|V|-2}\otimes\bra{x}(U_{f}^{\dagger}\otimes I_{n})\textsc{SWAP}(U_{f}\otimes I_{n})\ket{0}^{\otimes n+|V|-2}\otimes\ket{y}$ (39)
[…]
$=$ $\sum_{z,z^{\prime}}c^{\ast}_{z}c_{z^{\prime}}f(z)f(z^{\prime})\braket{z}{y}\braket{x}{z^{\prime}}$ (41)
$=$ $\braket{x}{\psi_{f}}\braket{\psi_{f}}{y}.$ (42)
As another example, for the above $\mathrm{WFBDD}_{f}$, if we can implement the controlled unitaries of $V_{z}\in SU(2^{n})$ efficiently, we can construct a $(1,n+|V|-2,0)$-block encoding of a $2^{n}\times 2^{n}$ matrix
$A=\sum_{z\in\{0,1\}^{n}}|c_{z}|^{2}f(z)V_{z},$ (43)
by following the technique of the block encoding of Gram matrices (Lemma 48 of Gilyén et al. (2018)). Specifically, we can construct two unitaries $U_{L}$, and $U_{R}$ such that
$U_{L}:\ket{0^{n}}\otimes\ket{0}^{\otimes|V|-2}\otimes\ket{y}$ $\mapsto$ $\sum_{z\in\{0,1\}^{n}}c_{z}f(z)\ket{z}\otimes\ket{0}^{\otimes|V|-2}\otimes V(z)\ket{y},$ (44)
$U_{R}:\ket{0^{n}}\otimes\ket{0}^{\otimes|V|-2}\otimes\ket{x}$ $\mapsto$ $\sum_{z\in\{0,1\}^{n}}c_{z}f(z)\ket{z}\otimes\ket{0}^{\otimes|V|-2}\otimes\ket{x},$ (45)
and then $U_{R}^{\dagger}U_{L}$ is the block encoding of $A$.
The above two results can give non-trivial examples of block encodings of non-sparse matrices. For example, the block encodings of matrices derived from the $\mathrm{WFBDD}_{h}$ of which $h$ is the Boolean function of Eq. (12) are the cases, since the $\mathrm{WFBDD}_{h}$ represents a $n$-qubit quantum state with $d=O(2^{n})$ nonzero coefficients, as shown in Section III.
VIII Conclusion
In this paper, we proposed a quantum state preparation (QSP) algorithm when the classical description of an $n$-qubit quantum state can be represented by a weighted free binary decision diagram (${\rm WFBDD}_{f}$reitbar P et al. (1995) encoding a Boolean function $f:\{0,1\}^{n}\to\{0,1\}$ with $N$ nodes, and weighted edges. We analyzed the computational complexity of the proposed QSP algorithm. We constructed a quantum circuit preparing the $n$-qubit quantum state with $O(N)$ uses of single- or two-qubit gates, and $N-2$ ancillary qubits.
The efficiency of our QSP algorithm depends on the compression rate of WFBDDs. When $N=O(\mathrm{poly}(n))$, our algorithm gives an $O(\mathrm{poly}(n))$-sized QSP circuit with $\mathrm{poly}(n)$ ancillary qubits, which reduces the ancillary qubits needed in existing QSP algorithms for a general class of quantum states Zhang et al. (2022); Sun et al. (2023); Rosenthal (2021). Furthermore, we gave an example of $O(\mathrm{poly}(n))$-sized $\mathrm{WFBDD}_{f}$ representing an $n$-qubit quantum state with exponential nonzero coefficients. From these results, we contributed to propose a new subclass of efficiently preparable quantum states based on WFBDDs.
We gave three examples of QSPs which our FBDD-based QSP algorithm can resolve efficiently. First, we explored to prepare a uniform superposition over a Boolean function, i.e., $\sum_{x\in\{0,1\}}f(x)\ket{x}/\sqrt{|f|}$. We showed that there exists an $O(N)$-sized classical algorithm assigning appropriate edge-weights to convert a given $\mathrm{FBDD}_{f}$ to the $\mathrm{WFBDD}_{f}$ which describes the uniform superposition. Our algorithm exponentially improves the number of CNOT gates used by another BDD-based QSP Mozafari et al. (2020), which also aims to prepare a uniform superposition over a Boolean function.
Second, to prepare another superposition state encoding a Boolean function, when the Boolean function is represented by a ${\rm FBDD}_{f}$ with $N$ nodes, we modified our QSP algorithm to construct a controlled phase operation |$\ket{x}\mapsto e^{i\theta f(x)}\ket{x}$ using at most $2N$ Toffoli gates, and $N-1$ ancillary qubits. This result provides an example of the instantiation of a phase oracle of a Boolean function, and from which we can prepare the quantum state $\sum_{x\in\{0,1\}}e^{i\theta f(x)}\ket{x}$ trivially.
Third, we give one example of ${\rm WFBDD}_{f}$ with $O(n^{2})$ nodes describing an $n$-qubit quantum state which cannot be prepared efficiently by the black-box QSP Bausch (2022) with respect to the number of amplitude amplifications. Thus, this example shows that the QSP algorithms are complementary to each other.
Furthermore, we refereed that our QSP algorithm can be applied to the block encodings introduced in Ref Gilyén et al. (2019). Our research is significant in that it shows a different class of the preparable quantum states characterized by weighted graphs encoding Boolean functions than those of previous works Zhang et al. (2022); Sun et al. (2023); Rosenthal (2021). As a future work, it is interesting to explore the problems with respect to Boolean functions encoded by FBDDs, and their solutions through the quantum state preparations.
Acknowledgements.
H.Y. was supported by JST PRESTO Grant Number JPMJPR201A, JPMJPR23FC, JSPS KAKENHI Grant Number JP23K19970, and MEXT Quantum Leap Flagship Program (MEXT QLEAP) JPMXS0118069605, JPMXS0120351339. M.M. was supported by MEXT QLEAP Q-LEAP JPMXS0118069605, and KAKENHI Grant No. 21H03394.
Appendix A Quantum Random Access Memory (QRAM)
To access the classical information in the BDD data structure quantumly, we consider a quantum random access memory (QRAM) that stores quantum-accessible classical bits. In the implicit model of QRAM in Matteo et al. (2019), locations in memory are addressed by $n$-bit strings $x_{1}x_{2}\cdots x_{n}$, and are queried by inputting the associated state ket to a circuit:
[…]
where $b_{x_{1}x_{2}\cdots x_{n}}\in\{0,1\}$ is the stored value. Formally, one can realize the QRAM by implementing a sequence of mixed-polarity multi-controlled Toffolis (MPMCTs) conditioned on the control bits representing the memory address of a $1$. There are two known decompositions of quantum circuits into the $1$-, and $2$-qubit operations of the Clifford+$T$ gate set. One is a large-depth, small-width circuit. The other is a small-depth, large-width circuit Matteo et al. (2019).
Let $2^{q}$ be the number of $1$s in a database. To derive the small-depth, large-width circuit, first, the address bits input to the first registers of qubits are copied to the other ancilla registers of qubits by using a log-depth cascade of CNOTs. Second, each register performs an MPMCT whose target is one of the additional $2^{q}$ qubits. These additional qubits state is prepared in a superposition over even-parity states. Finally, compute the parity, copy it to an additional, copy back, and then uncompute the parity, and the address fanout (See Fig. 4 in Matteo et al. (2019)). From Matteo et al. (2019), the required number of qubits (the width of the circuit) is $O(n2^{q})$, and the required depth of the circuit scales linearly in both $n$ and $q$.